OpenStudy (anonymous):

The equation of a line is shown below: y = fraction reading 1/5x - 9 What is the equation of a line which is perpendicular to this line and passes through (1, 58)? y = 5x - 63 y = -5x + 63 y = 9x - 63 y = 9x + 63 @freckles

OpenStudy (anonymous):

@nincompoop

OpenStudy (freckles):

y=1/5*x-9 what is the slope of this line?

OpenStudy (anonymous):

um can i have a hiny

OpenStudy (anonymous):

hint

OpenStudy (freckles):

y=mx+b has slope m and y-intercept b what I'm asking for is m in this case since I'm looking for slope

OpenStudy (freckles):

what is the number in front of x in y=1/5* x-9

OpenStudy (anonymous):

-9

OpenStudy (freckles):